OpenStudy (anonymous):

The lesson is evaluating limits Limit of x / x +1 as X approaches -1 1/0 = INFINITY OR NEGATIVE INFINITY ?

OpenStudy (hba):

\[\lim_{x \rightarrow -1}\frac{ x }{ x+1 }\] Apply limit

OpenStudy (hba):

@Reyjuanx10

OpenStudy (anonymous):

-1 / -1 +1 = -1/0 my question is what's the answer is it infinity or negative infinity

OpenStudy (hba):

Well it would be negative infinity :)

OpenStudy (hba):

\[\frac{ -1 }{ 0 }=-\infty \]